If scores are normally distributed with the mean of 30 and standard deviation of 5. what percent score is.


1
Expert's answer
2022-04-29T14:24:34-0400

(a)

"P(X>30)=0.5"

(b)

"P(X>37)=1-P(Z\le\dfrac{37-30}{5})=0.0808"

(c)

"P(28<X<34)=P(Z<\dfrac{34-30}{5})"

"-P(Z\le\dfrac{28-30}{5})\approx0.4436"